### v1

# Investigating the Impact of Trust in Multi-Human Multi-Robot Task Allocation

Abstract
Trust is essential in human-robot collaboration. Even more so in multi-human multi-robot teams where trust is vital to ensure teaming cohesion in complex operational environments. Yet, at the moment, trust is rarely considered a factor during task allocation and reallocation in algorithms used in multi-human, multi-robot collaboration contexts. Prior work on trust in single-human-robot interaction has identified that including trust as a parameter in human-robot interaction significantly improves both performance outcomes and human experience with robotic systems. However, very little research has explored the impact of trust in multi-human multi-robot collaboration, specifically in the context of task allocation. In this paper, we introduce a new trust model, the Expectation Comparison Trust (ECT) model, and employ it with three trust models from prior work and a baseline no-trust model to investigate the impact of trust on task allocation outcomes in multi-human multi-robot collaboration. Our experiment involved different team configurations, including 2 humans, 2 robots, 5 humans, 5 robots, and 10 humans, 10 robots. Results showed that using trust-based models generally led to better task allocation outcomes in larger teams (10 humans and 10 robots) than in smaller teams. We discuss the implications of our findings and provide recommendations for future work on integrating trust as a variable for task allocation in multi-human, multi-robot collaboration.
I Introduction
Trust is crucial in several human-robot collaboration (HRC) tasks, such as search and rescue operations [1, 2], disaster response [3, 4], industrial collaboration [5, 6, 7, 8], and other contexts requiring multi-agent robot collaboration [9, 10, 11, 12, 13, 14, 15]. To capture the different dimensions of trust in HRC, researchers have provided different definitions of trust, including characterizing it as the belief that a human or robot team member will perform as expected in a given task scenario [16]. The guarantee that an agent possesses enough capacity to complete a task in uncertain and dynamic environments [17]. The expectation that agents collaborating on a task or project will not harm each other [18]. Other researchers have also characterized it as the belief that collaborating agents will be willing to act on the information and recommendations given to them to enable them to complete a task [16]. In this research, we align with the definition that posits that trust is the expectation that a team member has the required and expected capacity to complete a task.
Figure 1: A conceptual illustration of the proposed Expectation Comparison Trust (ECT) Model.
Prior work by researchers has adopted different computational strategies for modeling trust, including using probabilistic, non-probabilistic, and graph-based approaches to capture and model the different dynamic and behavioral aspects of trust. However, although several research studies have explored trust in single-HRC, very limited work has explored trust in the context of multi-human multi-robot (MH-MR) collaboration [19, 20, 21, 22]. Considering the increasing proliferation of larger, more diverse MH-MR teams and collaborative activities, it is crucial to investigate how different trust models perform in complex, multi-agent task allocation scenarios and their value in enhancing task allocation outcomes in MH-MR collaboration that involve complex and sensitive tasks.
In this research, we investigated the impact of trust on MH-MR task allocation performance outcomes. Our investigation involved five trust models, including a no-trust baseline model, three models from prior work, Monir et al.[23], the Xu and Dudek Model[24], the Guo and Yang Model [25], and the Expectation Confirmation Trust (ECT) Model we introduce in this research. Our ECT Model draws inspiration from the expectation confirmation theory [26] and applies its core concept for modeling trust in MH-MR contexts. Specifically, the ECT model, as illustrated in Fig. 1, posits that trust is the interplay between the level of divergence in prior expectations of performance and actual performance results, focusing especially on human expectations of robot performance. By conducting this investigation, we seek to highlight the importance of trust in MH-MR contexts and provide insights into the strengths and limitations of different trust modeling approaches in MH-MR contexts. Our research question included: How do different trust models affect task allocation in teams of varying size and composition like 2 humans and 2 robots (2H-2R), 5 humans and 5 robots (5H-5R), and 10 humans and 10 robots (10H-10R)?
Through our research, we make the following contributions: 1) We propose the ECT model to provide robotics researchers and practitioners with the ability to model trust in MH-MR task allocation; 2) We conduct an experiment to investigate the impact of introducing trust in MH-MR teams and found that trust improved the performance of higher team configurations; and 3) We further show that there is no monolithic trust model, as other models performed well in the experiment. This finding shows that there is no monolithic trust model, hence the model suitable for a project depends on various factors, including the task objectives, team composition, and other related factors.
The rest of the paper is structured as follows: Section II provides a detailed background on the evolution of trust models in HRI. Section III describes our proposed trust methodology, the simulation environment, and the case study setup of our experimental design. Section IV presents the results of our comparative analysis, followed by a discussion of the findings and their implications in Section V. Finally, Section VI concludes the paper and outlines directions for future research.
II Related Works
II-A Evolution of Trust Models in Human-Robot Collaboration
Over the past decade, several researchers have explored different approaches for introducing computational approximates of trust into HRC systems, interfaces, and algorithms. The overall goal of this research area is to computationally capture and model the different aspects of human attitudes, behaviors, and heuristics related to trust in ways that can be used to improve performance and safety in HRC. The dimensions of prior work in this area have explored a mixture of techniques for capturing, quantifying, or predicting different trust-approximate variables.
Guo et al. [27] highlighted that trust-enabled human-robot teaming often relies on using trust models to operationalize and transform different ideals of trust into computational approaches that support human-robot teaming. They introduced the reverse psychology trust-behavior model and compared its performance against prior existing models to investigate its performance. Their work further proposed a trust-seeking reward function that balances team performance and trust preferences. Chen et al. [28] proposed trust-POMDP as a means of introducing trust variables into human-robot interaction in ways that improve performance. They tested the performance of their approach using simulation and real-life experiments and found that their approach of factoring in trust improves human-robot team performance. Yemini et al. [29] introduced a mathematical framework for determining teaming consensus, especially where there are malicious agents. They used stochastic trust values for probabilistic measures of trustworthiness between agents in a distributed consensus system. Findings from their research revealed that trust-based approaches played a significant role in enhancing and improving the security and reliability of distributed systems.
Other researchers have also explored trust using different models, including trust-aware motion planning for HRC systems [30], examining the impact of using the Theory of Mind for modeling trust-aware robot policy for HRC [31], investigating the conditions under which robots can trust human wayfinding information for navigational guidance [32], examining approaches for eliciting trust levels during live interaction in HRC [33], and highlighting areas where trust plays a crucial role in robot proficiency by foregrounding to robots the areas where their capabilities are better matched with a given task and vice versa [34]. Relevant here is that although several prior works have explored trust in HRC, very few research have explored trust in the context of complex and sometimes heterogenous MH-MR teams.
II-B Factors Influencing Trust Dynamics
Trust models often seek to quantify the complex interplay of psychological, social, and environmental factors present in HRC. However, researchers have an implicit understanding of the difficulty of achieving a complete model of trust that captures all the rich variables at play in HRC. As a result, prior work by researchers has focused on different trust elements that are peculiar and at play in trust formation and maintenance in a given context of HRC.
Hudspeth et al. [35] conducted a study to investigate the impact of interfaces on human-robot trust collaboration in physical environments. Findings from their study revealed that interfaces and feedback provided through them significantly impact levels of trust in HRC. Luebbers et al. [36] highlighted the role of human memory recall of a robot’s previous performance on trust in HRC, highlighting that a last memory of a robot failing a task leads to increased distrust of humans on robot competence. They conducted an experiment with 53 participants to test this assertion further. Findings from their study validated concerns about the impact of recency bias on human trust in robot competence. Hald and Rehm [37] used several questionnaires, including human-robot trust questionnaires, to elicit information about ways of improving the usability of HRC systems. Findings from their research using such trust questionnaires in systems led to increased usability and high trust in system performance. Rehm et al. [38] highlighted the ways in which trust plays a role in human-robot interaction in social contexts, especially with respect to users relying on or rejecting the assistance of robot-assisted shopping. Findings from their research showed that positive rejection increases user trust in the system.
Furthermore, Wang et al. [39] investigated the value of using trust for role arbitration in situations where control is alternated between humans and robots in ways that optimize team performance. They introduced a quantitative approach for modeling trust and self-confidence in HRC. Findings from their research revealed that their approach delivered superior results in HRC in achieving a task. Xu et al. [40] also investigated the approaches for enhancing trust recognition in HRC using precise approaches. They introduced an EEG-based approach that includes an EEG Vision Transformer and EEGTrust dataset for detecting and measuring levels of human trust during HRC. Findings from their research revealed that their approach produced significant results that outscored other baseline approaches. They further made their code and dataset available to facilitate this genre of research.
III Methods
III-A Trust Models
We employed different trust models for our experiment, including the no-trust baseline model, Monir et al. [23], Xu and Dudek Model [24], Guo and Yang Model [25], and the Expectation Confirmation Model (ECM), which we contribute in through this research. Overall, our objective was to examine the impact of the different trust models on task allocation in the context of MH-MR. We briefly describe each model below.
III-A1 No-trust baseline model
We employed the no-trust baseline model to investigate the outcome of not including trust as a parameter for task allocation in MH-MR contexts and to compare its performance to other trust-enabled models.
III-A2 Monir Trust Model
Monir et al. [23] treated trust as a step function that relies on the performance of the human and robot performance to measure trust levels. They categorized trust into three regions, including unpredictable, predictable, and faithful regions. They highlighted that the human-robot interaction starts with the robot in the dependable region and should ideally end at the faithful region if the interaction goes well. They provided the following equation as a means of modeling their trust concept:
$T(t)=\begin{cases}0,&\text{for }R_{p}(t)<f_{P}\\ \epsilon,&\text{for }f_{P}\leq R_{p}(t)<f_{D}\\ \min(1,\epsilon+\tanh(c\Delta P)),&\text{for }f_{D}\leq R_{p}(t)<f_{F}\\ 1,&\text{for }R_{p}(t)\geq f_{F}\end{cases}$ (1)
where $T$ is the trust level, $t$ is time, $R_{p}(t)$ is the robot’s performance at time $t$, $f_{P}$, $f_{D}$, and $f_{F}$ are thresholds for predictability, dependability, and faith, respectively, $c$ and $\epsilon$ are constants based on the person’s preferences, and $\Delta P=R_{p}(t)-f_{D}$.
III-A3 Xu and Dudek Trust Model
The Xu and Dudek Model [24] used a Dynamic Bayesian Network to model and infer trust in HRC. The model captures different human-robot interaction elements, including human interventions, robot performance, and the overall trust change reports during task completion. The core idea of their model is captured as follows:
$\begin{split}P(t_{k},t_{k-1},p_{k},p_{k-1})&\approx N(t_{k};t_{k-1}+\omega_{tb}+\omega_{tp}p_{k}\\ &\quad+\omega_{td}(p_{k}-p_{k-1}),\sigma_{t})\end{split}$ (2)
$O(t_{k}=1,t_{k-1},i_{k},e_{k}):=S(\omega_{ib}+\omega_{it}t_{k}+\omega_{id}1t_{k}+\omega_{ie}e_{k})$ (3)
$O_{c}(t_{k},t_{k-1},c_{k}):=\text{Prob}(c_{k}|t_{k},t_{k-1})$ (4)
where $N$ denotes a Gaussian distribution, $S(x)$ is the sigmoid function, $\omega_{tb}$, $\omega_{tp}$, and $\omega_{td}$ are weights related to trust, $p_{k}$ is the robot’s performance, $t_{k}$ is the trust state at time $k$, and $i_{k}$ and $e_{k}$ denote human intervention and task change, respectively.
III-A4 Guo and Yang Trust Model
The Guo and Yang model [25] is a personalized trust prediction model that uses a Beta distribution and Bayesian inference to measure trust in HRC. Their model appreciates that trust is not static but evolves as the interaction between the human and robot progresses. They validated their model by comparing it to other trust models that came before it. The core formulation of their model is as follows:
$\alpha_{i}=\begin{cases}\alpha_{i-1}+w_{s},&\text{if }p_{i}=1\\ \alpha_{i-1},&\text{if }p_{i}=0\end{cases}$ (5)
$\beta_{i}=\begin{cases}\beta_{i-1}+w_{f},&\text{if }p_{i}=1\\ \beta_{i-1},&\text{if }p_{i}=0\end{cases}$ (6)
$\hat{t}_{i}=\frac{\alpha_{i}}{\alpha_{i}+\beta_{i}}$ (7)
where $\alpha_{i}$ and $\beta_{i}$ are parameters of the Beta distribution, $p_{i}$ is the performance indicator (1 for success, 0 for failure), $w_{s}$ and $w_{f}$ are weights for positive and negative experiences, and $\hat{t}_{i}$ is the predicted trust at the completion of the $i$-th task.
III-A5 Expectation Confirmation Trust (ECT) Model
Figure 2: Illustration of the core process of our ECT Model.
The ECT Model is built from the Expectation Confirmation Theory, a technique from psychology which defines satisfaction as the interplay between expectation, performance, and disconfirmation [26]. In our context, the ECT model is built on the assumption that trust in the context of HRC is based on the difference between expectations of robot performance before a task and the actual performance of a robot following the completion of a task.
The ECT model addresses a critical gap by introducing trust models for MH-MR teams and explicitly incorporating the dynamic nature of human expectations in collaborative scenarios. Unlike previous single-human robot trust models that often treat trust as a static or slowly evolving parameter, the ECT model recognizes that human trust in robotic teammates is continuously shaped by the interplay between prior expectations and observed performance. This viewpoint is crucial in complex, high-stakes environments where team compositions and task demands change rapidly. By grounding the model in Expectation Confirmation Theory [26], an established framework from consumer psychology, we leverage insights from human decision-making processes that have been extensively studied and adapted in different contexts [41, 42]. This interdisciplinary approach allows us to capture nuanced aspects of trust formation and maintenance in new ways.
The ECT model includes several components as illustrated in Fig. 2, including initial trust, which is often based on the assumption of a robot’s capability and perception; performance evaluation, which is based on the actual performance of a robot on a task; trust updating mechanisms that seek to match initial expectations to performance, and team-trust aggregation function that seeks to determine the overall team trust on a robot capability as a means of measuring and determining the final trust score for task allocation. Our model outline is presented in detail in Table I.
TABLE I: Framework for ECT model for multi-human, multi-robot (MH-MR) systems. This framework addresses human trust in robot performance during MH-MR collaboration.
             Concept                       Definition                       Formula
[…]
Performance Comparison The robot’s actual performance is compared to the expected performance to determine the degree of expectation confirmation. $\Delta_{ir}=P_{ir}-E_{i-1r}$
Interpretation: $\Delta_{ir}>0$: Positive confirmation (robot exceeds expectations). $\Delta_{ir}\leq 0$: Negative or neutral confirmation.
Trust Update Function Trust in robot $r$ is updated based on the performance comparison ($\Delta_{ir}$) using a non-linear function to avoid drastic trust shifts due to isolated events. $T_{ir}=T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
where $\alpha$ is the learning rate, determining how quickly trust adjusts, $\beta$ is the scaling factor that controls the influence of $\Delta_{ir}$ on trust change, and the hyperbolic tangent function ($\tanh$) ensures changes in trust are gradual and bounded.
Trust Decay Trust diminishes over time in the absence of interactions, accounting for the fading impact of past experiences. $T_{ir}=(1-\gamma)\times T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
where $\gamma$ is the decay rate, indicating how quickly trust diminishes without recent interactions, and $T_{ir}=\min(\max(T_{ir},0),1).$
Trust Facets Trust is multi-dimensional, incorporating aspects like competence, ability, and dependability. $T_{ir}=w_{1}T_{\text{com},ir}+w_{2}T_{\text{abi},ir}+w_{3}T_{\text{dep},ir}$
where $w_{1},w_{2}$, and $w_{3}$ are weights reflecting the relative importance of each trust component.
[…]
Allocation Criteria: High Trust: Robots with higher trust levels receive more complex tasks. Low Trust: Robots with lower trust are assigned simpler tasks or monitored more closely.
Team Trust Aggregation Aggregates individual trust assessments to derive a collective trust measure for each robot. $T_{\text{team}r}(t)=\frac{1}{m}\sum_{i=1}^{m}T_{ir}(t)$
which provides an overall trust level that reflects the team’s collective confidence in each robot.
[…]
III-B Simulation
We implemented a simulation environment to evaluate the performance of various trust models in MH-MR task allocation. The simulation employed RL techniques to model agent behavior and decision-making processes. The simulation environment is represented by a 2D grid world of size 100 $\times$ 100 meters. A total of 10 Points of Interest (POIs) are randomly distributed across the map for each simulation run. The environment is defined as:
$E={G,P,T}$ (8)
where $G$ represents the grid world, $P$ is the set of POIs, and $T$ denotes the terrain type. Each POI $p_{i}\in P$ is associated with an attribute $a_{i}\in{{}^{\prime}survivor^{\prime},^{\prime}hazard^{\prime},^{\prime}resource^{\prime}}$, randomly assigned during initialization. Two types of agents are modeled: humans ($H$) and robots ($R$). Robots are further categorized into Unmanned Aerial Vehicles (UAVs) and Unmanned Ground Vehicles (UGVs). Each agent type $j$ has specific attributes:
$A_{j}={v_{j},s_{j}}$ (9)
where $v_{j}$ is the speed (in m/s) and $s_{j}$ is the sensing range (in meters).
Humans: $v_{H}=1.0$ m/s, $s_{H}=10$ m; UAVs: $v_{UAV}=2.0$ m/s, $s_{UAV}=20$ m; UGVs: $v_{UGV}=1.5$ m/s, $s_{UGV}=15$ m. These speeds are chosen to reflect realistic movement capabilities in SAR. The simulation progresses in discrete time steps, where each step represents 1 second of real-time. We implemented and compared five trust models: 1) NoTrust: A baseline model with no trust updates, 2) Monir: Based on Monir et al.’s work, using discrete trust regions, 3) XuDudek: Implementing Xu and Dudek’s continuous trust update mechanism, 4) GuoYang: Utilizing Guo and Yang’s Beta distribution-based trust model, and 5) ECT: Our proposed ECT model incorporates task complexity and robot capability. Tasks are allocated using a Q-learning approach [43]. The state space $S$ is defined by the agent’s distance to the nearest POI and its current trust level. The action space $A$ consists of three possible actions: move to POI, perform task, and wait. The Q-learning update rule is:
$Q(s,a)\leftarrow Q(s,a)+\alpha[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)]$ (10)
where $\alpha=0.1$ is the learning rate, $\gamma=0.9$ is the discount factor, and $r$ is the reward. The exploration rate $\epsilon$ is set to 0.1.
We conducted a series of experiments to evaluate the performance of different trust models across various team compositions and environmental conditions, including A) Small team (2 Humans and 2 Robots; 2H-2R) in flat terrain, B) Medium team (5 Humans and 5 Robots; 5H-5R) in rough terrain, and C) Large team (10 Humans and 10 Robots; 10H-10R) in obstacle-dense terrain. The environments types (e.g., flat terrain) represent levels of task complexity. We used two primary metrics to evaluate performance, including:
  • Task Success Rate (TSR): The proportion of successfully completed tasks.
    $TSR=\frac{\text{Number of completed tasks}}{\text{Total number of tasks}}$ (11)
  • Average Completion Time (ACT): The average time taken to complete tasks.
    $ACT=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (12)
where $t_{i}$ is the completion time of task $i$, and $n$ is the total number of completed tasks.
For each combination of trust model and scenario, we conducted 100 independent simulation runs to ensure statistical significance. Each run lasted for a maximum of 500 time steps or until all tasks were completed. We used the trust questionnaire from Yagoda and Gillian [44] as our theoretical framework for defining trust parameters, specifically focusing on measuring function and performance-related trust parameters, which include competence, ability, and dependability. The simulation logged key metrics at each step, including agent positions, trust levels, task completion times, and task success rates. This data allowed for a comparison of the trust models in terms of their impact on team performance and collaboration in a multi-agent SAR environment.
IV Results and Analysis
(a) Success rate
(b) Average of completion time
Figure 4: Results of the Success Rate and Average Completion Time of the Different Trust Models in team compositions. The star symbol indicates the best performance in each result, the highest success rate, and the shortest completion time. Zoom in for better readability.
Results from our evaluation of the five trust models, including the No Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across three team compositions (e.g., 2H-2R, 5H-5R, and 10H-10R), revealed close results in both task success rates and average task completion times. In small team configurations (2H-2R), we observed that all trust models performed comparably in terms of success rates. ECT, NoTrust, and XuDudek all achieved the highest success rate of 0.35 ± 0.08, while Monir and GuoYang followed closely at 0.33 ± 0.08. Regarding completion times, XuDudek was the fastest at 49.28 ± 16.72 seconds, with ECT following closely at 49.66 ± 14.43 seconds. Notably, ECT demonstrated the lowest standard deviation in completion time, indicating more consistent performance even in small-scale operations. As we moved to medium-sized teams (5H-5R), the performance differences between trust models became more pronounced. Monir took the lead in success rate at 0.63 ± 0.10, with ECT following closely at 0.62 ± 0.10. The other models (NoTrust, XuDudek, and GuoYang) tied at 0.60, albeit with slightly higher standard deviations. Interestingly, the completion time results for this team size revealed a different pattern. Monir was the fastest at 48.42 ± 11.73 seconds, while ECT had the slowest mean completion time at 52.25 ± 11.25 seconds. This suggests a potential trade-off in ECT between maintaining a high success rate and optimizing completion time in moderately complex scenarios.
The most notable results emerged in the large team configuration (10H-10R), where ECT demonstrated its strengths. In this complex scenario, ECT tied with GuoYang for the highest success rate at 0.82 ± 0.10, outperforming the other models. Moreover, ECT achieved the fastest mean completion time of 48.99 ± 9.13 seconds, significantly quicker than the next fastest model, Monir, at 49.14 ± 10.08 seconds. Also, the ECT model exhibited the lowest standard deviations in both success rate and completion time, indicating the most consistent and reliable performance in large-scale MH-MR contexts. An important observation from these results is the scalability demonstrated by the ECT model as its success rate improved from 0.35 in small teams to 0.82 in large teams, representing one of the most significant improvements among all models. This scalability suggests that the ECT model’s trust dynamics become effective as team complexity grows, a critical factor for real-world applications where scalability is often paramount. In the 10H-10R configuration, the ECT model outperformed other models in terms of consistency, demonstrating the lowest standard deviations in both success rate (±0.10) and completion time (±9.13s).
V Discussion
Non-Linearity in Team Size and Performance. Results from our experiment showed a non-linear relationship between team size and performance across the different trust models. This non-linearity aligns with findings from prior work [45] which showed that large autonomous systems often exhibit non-linear behaviors due to the complex interactions between individual robots. The superior performance of trust-based models, particularly the ECT model, in larger team configurations (10H-10R) suggests a potential ’scale threshold’ where trust becomes a critical factor in team efficiency and optimal performance. The findings from our research open up new avenues for research into the optimal team sizing and threshold where trust measurement becomes crucial for MH-MR collaborations.
Trust as a Computational Resource. The ECT’s model’s performance, especially in larger teams introduces a novel perspective of trust as a form of distributed computational resource that leverages the perception of team members vis-a-vis robot performance. Unlike traditional views of trust as a purely psychological construct, our model and the results from the experiment suggest that trust functions as a form of collective intelligence that scales with team size and requires active participation from all team members for an objective outcome. This conceptualization of ’trust as a form of computation’ could help researchers reimagine how we design and implement trust models in robotic systems, potentially leading to an easily scalable and adaptable approach for eliciting and computing trust for large MH-MR teams. The design of the ECT model implicitly means that as the team size increases, the computational power of trust also increases.
Task Allocation and Trust Heuristics in Large Teams. The performance discrepancies between trust-based and non-trust models were more pronounced in larger team configurations and raised important questions about the role of trust in mitigating complexity in MH-MR teaming and task allocation. As team size increases, the computational demands of task allocation grow as well. Our results suggest that trust might serve as an additional heuristic, in addition to other factors like fatigue and workload, to reduce this complexity by providing insight and allowing for more efficient task allocation that is based on empirical metrics in large-scale MH-MR collaboration settings. Specifically, the ECT model is a scalable trust model that is adaptive to different team configurations and sizes and performs relatively well regardless of team composition, showing its usefulness.
Advantages of the ECT Model. Another key benefit of the ECT model lies in its direct approach. While other trust models often rely on independent techniques like facial detection, behavioral cues, and other indirect and after the fact approaches, the ECT model relies on direct elicitation of trust scores from human team members for a task and continuously throughout the lifecycle of the task. This approach ensures that the trust levels of human teams on robot performance are continuously updated. Furthermore, the approach of team trust aggregation allows for congregating the trust scores of all team members in measuring trust levels. Through this approach, the trust rating of an individual team does not have an outsized influence on the overall trust final score generated by the team.
Study Limitations and Future Directions. While results from our study provided valuable insights on trust-based task allocation in MH-MR teams, it is important to also acknowledge the limitations of our study. First, we used simulated environments for our experiment. While this allows for controlled experiments that do not expose users to harm, it may not fully capture the complexities that are often present in real-world contexts. Furthermore, our task domain might not translate to other scenarios and domains. As per prior work, trust is not easy to model, and as a result, our direct question-based method might not capture all the relevant nuances required for measuring trust. Despite these limitations, our work provides a good foundation for future research in trust-based task allocation approaches for MH-MR teams.
VI Conclusion
In this paper, we investigated the impact of trust on task allocation in multi-human multi-robot teams. We employed five trust models, including N-Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across different types of teams that included 2H-2R, 5H-5R, and 10H-10R. Our simulation task involved an SAR activity. Findings from our research revealed that trust-based models outperformed the no-trust model most of the time in the 10H10R team configuration and that there was no noticeable difference in the lower configurations. These findings have implications for trust in the context of MH-MR task allocation. Future work could explore trust repair mechanisms for MH-MR teams where trust has been broken due to poor task performance.

### v2

@@
-Investigating the Impact of Trust in Multi-Human Multi-Robot Task Allocation
+Modeling and Evaluating Trust Dynamics in Multi-Human Multi-Robot Task Allocation
 Abstract
-Trust is essential in human-robot collaboration.
-Even more so in multi-human multi-robot teams where trust is vital to ensure teaming cohesion in complex operational environments.
-Yet, at the moment, trust is rarely considered a factor during task allocation and reallocation in algorithms used in multi-human, multi-robot collaboration contexts.
-Prior work on trust in single-human-robot interaction has identified that including trust as a parameter in human-robot interaction significantly improves both performance outcomes and human experience with robotic systems.
-However, very little research has explored the impact of trust in multi-human multi-robot collaboration, specifically in the context of task allocation.
-In this paper, we introduce a new trust model, the Expectation Comparison Trust (ECT) model, and employ it with three trust models from prior work and a baseline no-trust model to investigate the impact of trust on task allocation outcomes in multi-human multi-robot collaboration.
-Our experiment involved different team configurations, including 2 humans, 2 robots, 5 humans, 5 robots, and 10 humans, 10 robots.
-Results showed that using trust-based models generally led to better task allocation outcomes in larger teams (10 humans and 10 robots) than in smaller teams.
-We discuss the implications of our findings and provide recommendations for future work on integrating trust as a variable for task allocation in multi-human, multi-robot collaboration.
+Trust is essential in human-robot collaboration, particularly in multi-human, multi-robot (MH-MR) teams, where it plays a crucial role in maintaining team cohesion in complex operational environments.
+Despite its importance, trust is rarely incorporated into task allocation and reallocation algorithms for MH-MR collaboration.
+While prior research in single-human, single-robot interactions has shown that integrating trust significantly enhances both performance outcomes and user experience, its role in MH-MR task allocation remains underexplored.
+In this paper, we introduce the Expectation Confirmation Trust (ECT) Model, a novel framework for modeling trust dynamics in MH-MR teams.
+We evaluate the ECT model against five existing trust models and a no-trust baseline to assess its impact on task allocation outcomes across different team configurations (2H-2R, 5H-5R, and 10H-10R).
+Our results show that the ECT model improves task success rate, reduces mean completion time, and lowers task error rates.
+These findings highlight the complexities of trust-based task allocation in MH-MR teams.
+We discuss the implications of incorporating trust into task allocation algorithms and propose future research directions for adaptive trust mechanisms that balance efficiency and performance in dynamic, multi-agent environments.
 I Introduction
-Trust is crucial in several human-robot collaboration (HRC) tasks, such as search and rescue operations [1, 2], disaster response [3, 4], industrial collaboration [5, 6, 7, 8], and other contexts requiring multi-agent robot collaboration [9, 10, 11, 12, 13, 14, 15].
-To capture the different dimensions of trust in HRC, researchers have provided different definitions of trust, including characterizing it as the belief that a human or robot team member will perform as expected in a given task scenario [16].
-The guarantee that an agent possesses enough capacity to complete a task in uncertain and dynamic environments [17].
-The expectation that agents collaborating on a task or project will not harm each other [18].
-Other researchers have also characterized it as the belief that collaborating agents will be willing to act on the information and recommendations given to them to enable them to complete a task [16].
-In this research, we align with the definition that posits that trust is the expectation that a team member has the required and expected capacity to complete a task.
+Trust is crucial in any human–robot collaboration (HRC) system, as it underpins effective coordination, fosters shared situational awareness, and minimizes misjudgments or misuse of robotic capabilities [1].
+In HRC, trust has been defined in various ways, including as the belief that a human or robot teammate will perform as expected in a given task scenario [2] and as the assurance that an agent possesses sufficient capability to complete a task in uncertain and dynamic environments [3].
+A key challenge lies in effectively modeling trust and leveraging it for improved collaboration.
+Prior research has employed various computational strategies to model trust, including probabilistic [4], non-probabilistic [3], and graph-based approaches [5], each capturing different dynamic and behavioral aspects of trust.
+However, despite these efforts to model trust in dyadic HRC, limited research has focused on trust dynamics within multi-agent human–robot teams [6].
+Multi-human, multi-robot (MH-MR) teams [7, 8, 9, 10] are emerging as a promising paradigm for large-scale, high-stakes scenarios such as disaster response, environmental surveillance, and military operations.
+These teams leverage the complementary strengths of multiple humans and robots to enhance coordination, efficiency, and adaptability.
+Compared to traditional HRC, MH-MR interactions are more dynamic, occurring in diverse patterns such as human–human, human–robot, and hierarchical chain structures.
+As a result, traditional trust models may not be effective, as they often fail to account for the complex, multi-agent dependencies, shifting team compositions, and evolving situational contexts inherent in MH-MR collaboration.
 Figure 1: A conceptual illustration of the proposed Expectation Comparison Trust (ECT) Model.
-Prior work by researchers has adopted different computational strategies for modeling trust, including using probabilistic, non-probabilistic, and graph-based approaches to capture and model the different dynamic and behavioral aspects of trust.
-However, although several research studies have explored trust in single-HRC, very limited work has explored trust in the context of multi-human multi-robot (MH-MR) collaboration [19, 20, 21, 22].
-Considering the increasing proliferation of larger, more diverse MH-MR teams and collaborative activities, it is crucial to investigate how different trust models perform in complex, multi-agent task allocation scenarios and their value in enhancing task allocation outcomes in MH-MR collaboration that involve complex and sensitive tasks.
-In this research, we investigated the impact of trust on MH-MR task allocation performance outcomes.
-Our investigation involved five trust models, including a no-trust baseline model, three models from prior work, Monir et al.[23], the Xu and Dudek Model[24], the Guo and Yang Model [25], and the Expectation Confirmation Trust (ECT) Model we introduce in this research.
-Our ECT Model draws inspiration from the expectation confirmation theory [26] and applies its core concept for modeling trust in MH-MR contexts.
-Specifically, the ECT model, as illustrated in Fig.
-1, posits that trust is the interplay between the level of divergence in prior expectations of performance and actual performance results, focusing especially on human expectations of robot performance.
-By conducting this investigation, we seek to highlight the importance of trust in MH-MR contexts and provide insights into the strengths and limitations of different trust modeling approaches in MH-MR contexts.
-Our research question included: How do different trust models affect task allocation in teams of varying size and composition like 2 humans and 2 robots (2H-2R), 5 humans and 5 robots (5H-5R), and 10 humans and 10 robots (10H-10R)?
-Through our research, we make the following contributions: 1) We propose the ECT model to provide robotics researchers and practitioners with the ability to model trust in MH-MR task allocation; 2) We conduct an experiment to investigate the impact of introducing trust in MH-MR teams and found that trust improved the performance of higher team configurations; and 3) We further show that there is no monolithic trust model, as other models performed well in the experiment.
-This finding shows that there is no monolithic trust model, hence the model suitable for a project depends on various factors, including the task objectives, team composition, and other related factors.
-The rest of the paper is structured as follows: Section II provides a detailed background on the evolution of trust models in HRI.
-Section III describes our proposed trust methodology, the simulation environment, and the case study setup of our experimental design.
-Section IV presents the results of our comparative analysis, followed by a discussion of the findings and their implications in Section V.
-Finally, Section VI concludes the paper and outlines directions for future research.
+This research first examines how traditional HRC trust models [11, 5, 12, 13] perform in MH-MR teams.
+Through this investigation, we aim to emphasize the significance of trust in MH-MR teams and offer insights into the strengths and limitations of existing trust modeling approaches in these complex, collaborative settings.
+We then introduce the Expectation Confirmation Trust (ECT) Model, a novel trust framework inspired by expectation confirmation theory [14], and adapt its core principles to model trust in MH-MR environments.
+Specifically, as illustrated in Fig.
+1, the ECT model posits that trust emerges from the interplay between prior expectations of performance and actual performance outcomes, with a particular focus on how humans perceive robot performance.
+Its dynamic nature, along with the ability to pool trust across different team members, enhances overall team performance while facilitating cross-learning, where lower-performing agents can learn from the best-performing ones.
+By examining the divergence between expected and observed outcomes, the model provides a structured approach to trust calibration in dynamic, multi-agent interactions.
+The main contributions are summarized below:
+  • We investigate trust dynamics in the multi-human, multi-robot (MH-MR) context, an area that remains largely unexplored.
+Specifically, we analyze the performance of traditional HRC trust models in MH-MR settings.
+  • We introduce the ECT model, a more dynamic and efficient approach for trust modeling in MH-MR task allocation.
+The ECT model leverages expectation confirmation theory to better capture trust adaptation and decision-making in multi-agent teams.
+  • Through scalable experiments across varying team sizes: from 2 humans and 2 robots to 10 humans and 10 robots, we demonstrate the benefits of the ECT model in improving overall team performance.
+Specifically, we show how ECT facilitates a better trade-off between task efficiency, resource allocation, and adaptability in dynamic MH-MR environments.
 II Related Works
 II-A Evolution of Trust Models in Human-Robot Collaboration
-Over the past decade, several researchers have explored different approaches for introducing computational approximates of trust into HRC systems, interfaces, and algorithms.
-The overall goal of this research area is to computationally capture and model the different aspects of human attitudes, behaviors, and heuristics related to trust in ways that can be used to improve performance and safety in HRC.
-The dimensions of prior work in this area have explored a mixture of techniques for capturing, quantifying, or predicting different trust-approximate variables.
+Over the past decade, researchers have developed various computational approaches to approximate trust in various HRC scenarios, like industrial collaboration [15, 16] and assistive care [17, 18].
+The overarching goal of this research is to model human attitudes, behaviors, and heuristics related to trust in ways that enhance team performance, safety, and adaptability.
+Prior work in this area has explored diverse techniques to capture, quantify, and predict trust-related variables.
 Guo et al.
-[27] highlighted that trust-enabled human-robot teaming often relies on using trust models to operationalize and transform different ideals of trust into computational approaches that support human-robot teaming.
-They introduced the reverse psychology trust-behavior model and compared its performance against prior existing models to investigate its performance.
-Their work further proposed a trust-seeking reward function that balances team performance and trust preferences.
-Chen et al.
-[28] proposed trust-POMDP as a means of introducing trust variables into human-robot interaction in ways that improve performance.
-They tested the performance of their approach using simulation and real-life experiments and found that their approach of factoring in trust improves human-robot team performance.
+[19] emphasized that trust-enabled human-robot teaming relies on computational trust models to operationalize and transform abstract trust concepts into actionable mechanisms that enhance collaboration.
+They introduced a reverse psychology trust-behavior model and compared its effectiveness against existing models, proposing a trust-seeking reward function to balance team performance and trust preferences.
+Similarly, Chen et al.
+[20] introduced trust-POMDP, integrating trust variables into decision-making processes to optimize human-robot interaction.
+Their approach, validated through simulation and real-world experiments, demonstrated improved team performance when factoring trust into HRC systems.
+Other studies have explored trust modeling from different perspectives.
 Yemini et al.
-[29] introduced a mathematical framework for determining teaming consensus, especially where there are malicious agents.
-They used stochastic trust values for probabilistic measures of trustworthiness between agents in a distributed consensus system.
-Findings from their research revealed that trust-based approaches played a significant role in enhancing and improving the security and reliability of distributed systems.
-Other researchers have also explored trust using different models, including trust-aware motion planning for HRC systems [30], examining the impact of using the Theory of Mind for modeling trust-aware robot policy for HRC [31], investigating the conditions under which robots can trust human wayfinding information for navigational guidance [32], examining approaches for eliciting trust levels during live interaction in HRC [33], and highlighting areas where trust plays a crucial role in robot proficiency by foregrounding to robots the areas where their capabilities are better matched with a given task and vice versa [34].
-Relevant here is that although several prior works have explored trust in HRC, very few research have explored trust in the context of complex and sometimes heterogenous MH-MR teams.
+[21] developed a mathematical framework for consensus in multi-agent teams, particularly in environments with malicious agents, utilizing stochastic trust values for probabilistic trust assessment.
+Their findings highlighted the role of trust in enhancing security and reliability in distributed systems.
+Yu et al.
+[22] investigated trust-aware motion planning for HRC, while another study by the same authors [23] examined the application of Theory of Mind in trust-aware robot policy learning.
+Additionally, researchers have explored how trust influences wayfinding guidance, live trust elicitation in HRC [24], and adaptive trust-based task allocation where robots assess their capabilities against task requirements.
+While these studies have significantly advanced trust modeling in HRC, research on trust within complex, heterogeneous MH-MR teams remains scarce.
+Given the increasing prevalence of large-scale MH-MR deployments in high-stakes environments, there is a critical need to extend trust modeling frameworks beyond single-human, single-robot interactions to account for dynamic, multi-agent collaboration and adaptive trust calibration in MH-MR contexts.
 II-B Factors Influencing Trust Dynamics
@@
 Hudspeth et al.
-[35] conducted a study to investigate the impact of interfaces on human-robot trust collaboration in physical environments.
-Findings from their study revealed that interfaces and feedback provided through them significantly impact levels of trust in HRC.
+[25] found that interface design and feedback significantly influence trust levels in physical HRC.
 Luebbers et al.
-[36] highlighted the role of human memory recall of a robot’s previous performance on trust in HRC, highlighting that a last memory of a robot failing a task leads to increased distrust of humans on robot competence.
-They conducted an experiment with 53 participants to test this assertion further.
-Findings from their study validated concerns about the impact of recency bias on human trust in robot competence.
-Hald and Rehm [37] used several questionnaires, including human-robot trust questionnaires, to elicit information about ways of improving the usability of HRC systems.
-Findings from their research using such trust questionnaires in systems led to increased usability and high trust in system performance.
-Rehm et al.
-[38] highlighted the ways in which trust plays a role in human-robot interaction in social contexts, especially with respect to users relying on or rejecting the assistance of robot-assisted shopping.
-Findings from their research showed that positive rejection increases user trust in the system.
-Furthermore, Wang et al.
-[39] investigated the value of using trust for role arbitration in situations where control is alternated between humans and robots in ways that optimize team performance.
-They introduced a quantitative approach for modeling trust and self-confidence in HRC.
-Findings from their research revealed that their approach delivered superior results in HRC in achieving a task.
+[26] demonstrated that recent failures lead to increased distrust due to recency bias.
+Hald and Rehm [27] showed that integrating trust questionnaires improves usability and system trust, and found that positive rejection in robot-assisted shopping can increase trust.
+Beyond cognitive and social factors, Wang et al.
+[28] introduced a trust-driven role arbitration model to optimize control transitions between humans and robots.
 Xu et al.
-[40] also investigated the approaches for enhancing trust recognition in HRC using precise approaches.
-They introduced an EEG-based approach that includes an EEG Vision Transformer and EEGTrust dataset for detecting and measuring levels of human trust during HRC.
-Findings from their research revealed that their approach produced significant results that outscored other baseline approaches.
-They further made their code and dataset available to facilitate this genre of research.
-III Methods
+[29] proposed an EEG-based trust recognition framework, leveraging physiological signals for real-time trust assessment [30].
+While these studies provide valuable insights into trust modeling in single-human, single-robot interactions, they do not generalize well to MH-MR collaboration, where trust must be dynamically pooled, distributed, and updated across multiple agents.
+Unlike prior works that focus on individual trust quantification or role arbitration, our proposed ECT Model is inspired by expectation confirmation theory, which explicitly accounts for the discrepancy between expected and observed performance across multiple human and robot agents.
+Unlike static models that assume trust is an individual attribute, our model dynamically adjusts trust at the team level, enabling trust redistribution and cross-learning between high- and low-performing agents.
+III Methodology
 III-A Trust Models
-We employed different trust models for our experiment, including the no-trust baseline model, Monir et al.
-[23], Xu and Dudek Model [24], Guo and Yang Model [25], and the Expectation Confirmation Model (ECM), which we contribute in through this research.
+We consider different trust models including the no-trust baseline model, Monir et al.
+[11], the Xu and Dudek Model [5], the Guo and Yang Model [12]), the Ali et al.
+Model [13], the Guo et al.
+Model [6], and our Expectation Confirmation Trust model (ECT), which we contribute through this research.
 Overall, our objective was to examine the impact of the different trust models on task allocation in the context of MH-MR.
@@
 III-A1 No-trust baseline model
-We employed the no-trust baseline model to investigate the outcome of not including trust as a parameter for task allocation in MH-MR contexts and to compare its performance to other trust-enabled models.
+We employed the no-trust baseline model to investigate the outcome of not including trust as a parameter for task allocation in MH-MR contexts.
 III-A2 Monir Trust Model
@@
 where $\alpha_{i}$ and $\beta_{i}$ are parameters of the Beta distribution, $p_{i}$ is the performance indicator (1 for success, 0 for failure), $w_{s}$ and $w_{f}$ are weights for positive and negative experiences, and $\hat{t}_{i}$ is the predicted trust at the completion of the $i$-th task.
-III-A5 Expectation Confirmation Trust (ECT) Model
+III-A5 Ali et al. Trust Model
+Ali et al.
+introduce an artificial trust measure for an agent $a$ performing a task $\gamma$ at time $t$ by comparing the task’s requirements $\bar{\alpha}_{i}$ to the agent’s current belief bounds $\ell_{i}^{a}(t)$ and $u_{i}^{a}(t)$ for each capability dimension $i=1,\dots,n$.
+The resulting “one-shot” trust value $\tau_{a,\gamma}(t)$ is given by
+$\tau_{a,\gamma}(t)\;=\;\prod_{i=1}^{n}\psi\bigl{(}\bar{\alpha}_{i}\bigr{)},$ (8)
+where the function $\psi(\bar{\alpha}_{i})$ compares $\bar{\alpha}_{i}$ against the current lower and upper bounds:
+[…]
+Hence, if the task requirement $\bar{\alpha}_{i}$ falls below the agent’s believed lower bound for capability $i$, trust on that dimension is $1$; if it exceeds the agent’s believed upper bound, trust on that dimension is 00; and otherwise it scales linearly in between.
+After each interaction (success/failure), these bounds $\ell_{i}^{a}(t)$ and $u_{i}^{a}(t)$ can be updated accordingly (e.g., expanded upon successful performances or shrunk upon failure).
+III-A6 Guo et al. Trust Model
+Guo et al.
+define trust $t_{a,b,k}$ as a Beta-distributed random variable that evolves over discrete interaction steps $k$.
+For agent $a$ to trust agent (or robot) $b$, they maintain Beta parameters $(\alpha_{a,b,k},\beta_{a,b,k})$ and update them after each direct or indirect experience:
+$t_{a,b,k}\;\sim\;\text{Beta}\Bigl{(}\alpha_{a,b,k},\,\beta_{a,b,k}\Bigr{)}.$ (10)
+The parameters $\alpha_{a,b,k}$ and $\beta_{a,b,k}$ typically increase based on positive and negative experiences, respectively.
+Concretely, if $p^{k}_{b}$ indicates a success/failure term for the $k$-th interaction (e.g., $p^{k}_{b}=1$ for success, $=0$ for failure), Guo et al.
+add coefficients $s_{a,b}$ and $f_{a,b}$ to reflect how strongly each experience changes $\alpha$ or $\beta$:
+$\alpha_{a,b,k}\;=\;\alpha_{a,b,k-1}\;+\;s_{a,b}\,\bigl{(}\!\text{positive counts}\bigr{)},$ (11)
+$\beta_{a,b,k}\;=\;\beta_{a,b,k-1}\;+\;f_{a,b}\,\bigl{(}\!\text{negative counts}\bigr{)}.$ (12)
+Afterward, the trust value $t_{a,b,k}$ is taken as the mean (or another statistic) of the updated Beta distribution, e.g.
+$t_{a,b,k}\;=\;\frac{\alpha_{a,b,k}}{\alpha_{a,b,k}+\beta_{a,b,k}}.$
+More variations may also factor in indirect experiences or testimonies from other agents (i.e., other humans/robots) by incrementing $\alpha_{a}$ or $\beta_{a}$ based on second-hand observations.
+III-A7 Expectation Confirmation Trust (ECT) Model
 Figure 2: Illustration of the core process of our ECT Model.
@@
 This viewpoint is crucial in complex, high-stakes environments where team compositions and task demands change rapidly.
-By grounding the model in Expectation Confirmation Theory [26], an established framework from consumer psychology, we leverage insights from human decision-making processes that have been extensively studied and adapted in different contexts [41, 42].
-This interdisciplinary approach allows us to capture nuanced aspects of trust formation and maintenance in new ways.
-The ECT model includes several components as illustrated in Fig.
-2, including initial trust, which is often based on the assumption of a robot’s capability and perception; performance evaluation, which is based on the actual performance of a robot on a task; trust updating mechanisms that seek to match initial expectations to performance, and team-trust aggregation function that seeks to determine the overall team trust on a robot capability as a means of measuring and determining the final trust score for task allocation.
-Our model outline is presented in detail in Table I.
+By grounding the model in Expectation Confirmation Theory [14], we leverage insights from human decision-making processes that have been extensively studied and adapted in different contexts [31, 32].
+As illustrated in 3, the ECT model consists of several components: (1) initial trust, based on assumptions of a robot’s capability and perception; (2) performance evaluation, which assesses a robot’s actual task performance; and (3) trust updating mechanisms that align initial expectations with observed performance.
+Our model outline is presented in detail in Table I and described below.
+In this work, we implemented the ECT model by incorporating multi-faceted trust, adaptive learning rates, task-specific histories, and partial/conditional decay.
+Our overall goals were (i) to evaluate robot performance against evolving expectations, (ii) to update trust accordingly, and (iii) to decide if a given agent should attempt or decline a task under different contexts based on their score.
+Below, we detail each key component.
 TABLE I: Framework for ECT model for multi-human, multi-robot (MH-MR) systems. This framework addresses human trust in robot performance during MH-MR collaboration.
-             Concept                       Definition                       Formula
+             Concept                       Definition                       Formula / Note
 […]
+Initial Expectation ($E_{0r}$) Initial expectations for robot $r$ are set based on the task’s complexity, the environment, and the robot’s (initial) capabilities.
+Our model (ECT) allows skill to improve over time (cross-agent learning).
+$E_{0r}=g(\mathrm{complexity},\mathrm{robot\_skill})$ Our implementation example: $E=\mathrm{skill}\times(1.2-0.1\times(\mathrm{complexity}-1))$
+Performance Evaluation ($P_{ir}$) After the robot executes a task ($i$), its performance is evaluated using metrics such as completion time, accuracy, and adaptability.
+Successful completion can increase the robot’s skill levels in ECT, thus influencing future $E_{(i+1)r}$.
 […]
-[…]
-Performance Comparison The robot’s actual performance is compared to the expected performance to determine the degree of expectation confirmation. $\Delta_{ir}=P_{ir}-E_{i-1r}$
+Performance Comparison The robot’s actual performance is compared to the expected performance to determine the degree of expectation confirmation. $\Delta_{ir}=P_{ir}-E_{i-1,r}$
 Interpretation: $\Delta_{ir}>0$: Positive confirmation (robot exceeds expectations). $\Delta_{ir}\leq 0$: Negative or neutral confirmation.
-Trust Update Function Trust in robot $r$ is updated based on the performance comparison ($\Delta_{ir}$) using a non-linear function to avoid drastic trust shifts due to isolated events.
-$T_{ir}=T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
-where $\alpha$ is the learning rate, determining how quickly trust adjusts, $\beta$ is the scaling factor that controls the influence of $\Delta_{ir}$ on trust change, and the hyperbolic tangent function ($\tanh$) ensures changes in trust are gradual and bounded.
-Trust Decay Trust diminishes over time in the absence of interactions, accounting for the fading impact of past experiences.
-$T_{ir}=(1-\gamma)\times T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
-where $\gamma$ is the decay rate, indicating how quickly trust diminishes without recent interactions, and $T_{ir}=\min(\max(T_{ir},0),1).$
-Trust Facets Trust is multi-dimensional, incorporating aspects like competence, ability, and dependability. $T_{ir}=w_{1}T_{\text{com},ir}+w_{2}T_{\text{abi},ir}+w_{3}T_{\text{dep},ir}$
-where $w_{1},w_{2}$, and $w_{3}$ are weights reflecting the relative importance of each trust component.
-[…]
-Allocation Criteria: High Trust: Robots with higher trust levels receive more complex tasks. Low Trust: Robots with lower trust are assigned simpler tasks or monitored more closely.
-Team Trust Aggregation Aggregates individual trust assessments to derive a collective trust measure for each robot. $T_{\text{team}r}(t)=\frac{1}{m}\sum_{i=1}^{m}T_{ir}(t)$
-which provides an overall trust level that reflects the team’s collective confidence in each robot.
-[…]
+Trust Update Function Trust in robot $r$ is updated based on $\Delta_{ir}$ using a non-linear (tanh) function.
+Note: An adaptive learning rate $\alpha_{\mathrm{dyn}}$ that increases with consistent success and decreases otherwise. Partial or selective decay (smaller if recent performance is consistent).
+$T_{ir}=T_{i-1,r}+\alpha_{\mathrm{dyn}}(\text{consistency})\,\times\,\tanh\bigl{(}\beta\times\Delta_{ir}\bigr{)}$
+where $\alpha_{\mathrm{dyn}}$ depends on recent performance variance/success rates, $\beta$ controls the influence of $\Delta_{ir}$ on trust change, and $\tanh$ ensures changes are bounded.
+Trust Decay Trust diminishes if no recent successes, but selective decay reduces how much it fades when performance is consistent.
+$T_{ir}=\bigl{(}1-\gamma_{\mathrm{eff}}\bigr{)}\times T_{i-1,r}\;+\;\alpha_{\mathrm{dyn}}\,\tanh\bigl{(}\beta\times\Delta_{ir}\bigr{)},$
+where $\gamma_{\mathrm{eff}}$ is the effective decay rate modulated by recent performance;
+$T_{ir}=\min\bigl{(}\max(T_{ir},0),1\bigr{)}.$
+Trust Facets Trust is multi-dimensional, e.g., competence, reliability, and adaptability, each updated via the above logic.
+$T_{ir}=w_{1}\,T_{\mathrm{com},ir}\;+\;w_{2}\,T_{\mathrm{rel},ir}\;+\;w_{3}\,T_{\mathrm{ada},ir},$
+where $w_{1},w_{2},$ and $w_{3}$ reflect each facet’s importance.
+Task Allocation Decision Updated trust guides which tasks the robot attempts.
+Enhanced ECT also considers task-switching cost and distance/skill factors so that a more suitable agent may handle the task $A_{i+1,r}=f\bigl{(}T_{ir},\,\text{task priority},\,\text{team workload},\\[-2.15277pt] \text{switching cost},\,\text{distance factor}\bigr{)}$
+Allocation Criteria: High Trust: Robots with higher trust levels receive more complex tasks. Low Trust: Robots with lower trust handle simpler tasks.
+Multi-Faceted Trust
+Each agent $r$ maintains trust values across three distinct facets: competence, reliability, and adaptability. In vector form, we denote:
+$T_{r}(t)\;=\;\bigl{(}T_{\mathrm{comp},r}(t),~{}T_{\mathrm{rel},r}(t),~{}T_{\mathrm{ada},r}(t)\bigr{)}.$
+Each facet starts at an initial default trust (e.g.
+$0.7$).
+(In real-world this would be the human trust rating scores).
+When an agent attempts a task with complexity $c$ and obtains a performance outcome (reward) $P_{ir}\in[0,1]$, each trust facet is updated based on (i) how the actual performance compares with prior expectations, (ii) a non-linear $\tanh(\cdot)$ update, and (iii) partial decay that depends on recent consistency.
+Expectations & Task Histories
+For each task type (e.g.
+“hazard detection”) or attribute, the model tracks (a) a “task-specific trust” vector, (b) an expected performance $E_{i-1,r}$, and (c) a reward history $\{\,P_{1r},\dots,P_{i-1,r}\}$.
+This history is used to measure consistency, typically quantified by the variance or standard deviation of recent outcomes.
+A small variance (i.e.
+stable, repeated successes) raises the adaptive learning rate $\alpha_{\mathrm{dyn}}$ for that facet, while inconsistent or sporadic performance reduces it.
+Performance Comparison
+When an agent completes a task $i$, its actual performance $P_{ir}$ (a scalar reward) is compared to the prior expectation $E_{i-1,r}$. We define the “performance difference” as
+$\Delta_{ir}\;=\;P_{ir}\;-\;E_{i-1,r}.$
+If $\Delta_{ir}>0$, the robot is exceeding the expected baseline; if $\Delta_{ir}\leq 0$, performance is at or below expectation.
+Trust Update
+After the difference $\Delta_{ir}$ is computed, we perform a tanh-based update for each facet $f\in\{\mathrm{comp},\mathrm{rel},\mathrm{ada}\}$:
+$T_{f,r}(t)\;\leftarrow\;T_{f,r}(t^{-})\;+\;\alpha_{\mathrm{dyn}}\,\tanh\bigl{(}\beta\,\Delta_{ir}\bigr{)},$ (13)
+where $T_{f,r}(t^{-})$ is the facet’s trust just before the update, $\beta$ scales the influence of $\Delta_{ir}$, and $\alpha_{\mathrm{dyn}}$ is the adaptive learning rate, given by
+$\alpha_{\mathrm{dyn}}=\alpha_{\mathrm{base}}\times\Bigl{(}1+\kappa_{\mathrm{cons}}\times\mathbbm{1}\{\text{consistent success}\}\Bigr{)}\\ \times\bigl{(}1+\mathrm{varPenalty}\bigr{)}^{-1}.$ (14)
+Here, $\alpha_{\mathrm{base}}$ is a small constant (e.g.
+$0.05$), $\kappa_{\mathrm{cons}}>0$ is an added boost if the last few rewards were consistently above $0.7$, and $\mathrm{varPenalty}$ is proportional to the recent outcome variance.
+The $\tanh(\cdot)$ bounds the trust increment to avoid runaway growth.
+Selective Decay
+Finally, a selective decay factor is applied:
+$T_{f,r}(t)\;\leftarrow\;T_{f,r}(t)\times\bigl{(}1-\gamma_{\mathrm{eff}}\bigr{)},$ (15)
+where
+$\gamma_{\mathrm{eff}}\;=\;\gamma_{0}\,\times\,(1-\mathrm{consistency})\,\times\,\delta(\Delta_{ir}).$
+In our experiment $\gamma_{0}$ is a baseline decay rate (e.g.
+0.02), consistency is high if recent standard deviation is small, and $\delta(\Delta_{ir})\in\{0.5,1.0\}$ is lower for successful tasks ($\Delta_{ir}\geq 0$) and higher for failures.
+This ensures that if the agent performs well consistently, trust decays only minimally, but if failures occur or consistency is low, trust decays more.
+Task-Allocation Overrides
+Besides updating trust, the ECT model can override an agent’s normal Q-learning action whenever its trust or skill knowledge deems the task unsuitable. Specifically, a function
+$\mathrm{should\_attempt\_task}(r,\,\mathrm{task})$ (16)
+checks whether $(i)$ the agent’s competence/trust is sufficiently high, $(ii)$ no other agent is significantly better suited, and $(iii)$ the agent will not pay an excessive task-switching cost by abruptly changing task types.
+If these conditions fail, the agent may ignore the Q-learning policy and refuse the task, thereby deferring to a more suitable teammate.
+Cross-Agent Skill Learning
+The ECT model also increments the agent’s skill if it succeeds on a task, especially with higher complexity.
+For instance, success at a complex hazard task may provide a small skill bump, increasing future $E_{r}$ or the reliability facet.
+Over many interactions, the agent’s trust and skill can co-evolve, ensuring that consistent good performance not only raises trust but also improves the agent’s underlying capability parameter.
+Thus, the ECT model integrates:
+  1. 1.
+    multi-faceted trust (competence, reliability, adaptability),
+  2. 2.
+    adaptive trust updates (via $\tanh(\Delta)$ plus variable learning rates),
+  3. 3.
+    selective decay that depends on recent consistency and success/failure,
+  4. 4.
+    task-override logic that can bypass Q-learning decisions when trust or skill is insufficient (or if a better agent is available),
+  5. 5.
+    incremental skill increases upon successful task completion.
+These elements combined, allow trust to reflect both short-term performance outcomes and broader trends of reliability, so that the willingness of agents to tackle more complex tasks evolves over time in a team setting.
 III-B Simulation
@@
 The simulation progresses in discrete time steps, where each step represents 1 second of real-time.
-We implemented and compared five trust models: 1) NoTrust: A baseline model with no trust updates, 2) Monir: Based on Monir et al.’s work, using discrete trust regions, 3) XuDudek: Implementing Xu and Dudek’s continuous trust update mechanism, 4) GuoYang: Utilizing Guo and Yang’s Beta distribution-based trust model, and 5) ECT: Our proposed ECT model incorporates task complexity and robot capability.
-Tasks are allocated using a Q-learning approach [43].
-The state space $S$ is defined by the agent’s distance to the nearest POI and its current trust level.
-The action space $A$ consists of three possible actions: move to POI, perform task, and wait.
-The Q-learning update rule is:
-$Q(s,a)\leftarrow Q(s,a)+\alpha[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)]$ (10)
-where $\alpha=0.1$ is the learning rate, $\gamma=0.9$ is the discount factor, and $r$ is the reward. The exploration rate $\epsilon$ is set to 0.1.
+We implemented and compared seven trust models: 1) NoTrust: A baseline model with no trust updates, 2) Monir: Based on Monir et al.’s work, using discrete trust regions, 3) XuDudek: Implementing Xu and Dudek’s continuous trust update mechanism, 4) GuoYang: Utilizing Guo and Yang’s Beta distribution-based trust model, 5) Ali et al.
+Bounds-Based Trust Model, Guo et al.
+Beta-Distribution Trust Model, 7) ECT: Our proposed ECT model incorporates task complexity and robot capability.
+Tasks are allocated and executed via a Q-learning approach [33], where each agent chooses among three actions: move to the POI, perform the task, or wait.
+We define the state space
+$S=\bigl{(}\text{distance to nearest POI},\,\text{current trust level},\\ \text{POI complexity},\,\text{skill match}\bigr{)},$ (19)
+and the action space
+$A=\{\text{move to POI},\,\text{perform task},\,\text{wait}\}.$
+At each step, the agent observes the current state $s\in S$, takes an action $a\in A$, and receives a reward $r$ (e.g., positive for successful task completion, negative for failures or wasted movement).
+We use the standard Q-learning update rule:
+$Q(s,a)\;\leftarrow\;Q(s,a)\;+\;\alpha\Bigl{[}r+\gamma\,\max_{a^{\prime}}\,Q(s^{\prime},a^{\prime})\;-\;Q(s,a)\Bigr{]},$ (20)
+where $\alpha=0.1$ is the learning rate, $\gamma=0.9$ is the discount factor, and $s^{\prime}$ is the next state after executing action $a$.
+We employ an $\epsilon$-greedy policy with $\epsilon=0.1$ for exploration, meaning agents choose a random action with probability $0.1$ and the Q-learning action otherwise.
+Figure 3: Illustration of our virtual simulation environment containing different agents (UGVs, UAVs, and humans) and different points of interests.
 We conducted a series of experiments to evaluate the performance of different trust models across various team compositions and environmental conditions, including A) Small team (2 Humans and 2 Robots; 2H-2R) in flat terrain, B) Medium team (5 Humans and 5 Robots; 5H-5R) in rough terrain, and C) Large team (10 Humans and 10 Robots; 10H-10R) in obstacle-dense terrain.
-The environments types (e.g., flat terrain) represent levels of task complexity.
+The environment types (e.g., flat terrain) represent levels of task complexity.
 We used two primary metrics to evaluate performance, including:
   • Task Success Rate (TSR): The proportion of successfully completed tasks.
-    $TSR=\frac{\text{Number of completed tasks}}{\text{Total number of tasks}}$ (11)
   • Average Completion Time (ACT): The average time taken to complete tasks.
-    $ACT=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (12)
-where $t_{i}$ is the completion time of task $i$, and $n$ is the total number of completed tasks.
-For each combination of trust model and scenario, we conducted 100 independent simulation runs to ensure statistical significance.
+  • Agent Utilization: Tracks the percentage of simulation time during which agents are actively engaged in tasks rather than idle.
+Higher utilization indicates more efficient task allocation and less wasted time.
+  • Task Allocation Efficiency: Calculates the number of tasks completed per distance traveled, indicating how effectively agents are navigating to and selecting tasks.
+Higher values suggest smarter path planning and better task assignment.
+  • Error Rate: Measures the ratio of failed task attempts to total attempts.
+A lower error rate implies better trust calibration and decision-making, as agents avoid attempting tasks beyond their capabilities.
+  • Throughput: The number of tasks completed per unit time (or per time step). Higher throughput indicates that tasks are being completed more rapidly.
+  • Total Distance / Distance Per Task: The total distance traveled by all agents, as well as the average distance per completed task.
+These metrics capture navigation efficiency and resource expenditure.
+  • Collision Rate: The frequency of collisions among agents. Fewer collisions generally indicate better coordination and route planning.
+For each combination of trust model and scenario, we conducted 50 independent simulation runs to ensure statistical significance.
 Each run lasted for a maximum of 500 time steps or until all tasks were completed.
@@
 This data allowed for a comparison of the trust models in terms of their impact on team performance and collaboration in a multi-agent SAR environment.
+TABLE II: Performance metrics across trust models.
+Trust Model Team Mean SR Std SR Mean CT Std CT Simple TS Med TS Comp TS Error AU AE
+NoTrust 2H-2R 0.19 0.03 12.85 13.47 0.21 0.17 0.15 0.37 0.15 0.0868
+NoTrust 5H-5R 0.41 0.07 27.42 19.47 0.39 0.46 0.39 0.54 0.15 0.0701
+NoTrust 10H-10R 0.62 0.09 23.59 13.78 0.63 0.63 0.61 0.57 0.14 0.0530
+Monir 2H-2R 0.19 0.03 16.03 18.49 0.19 0.19 0.18 0.39 0.14 0.0848
+Monir 5H-5R 0.39 0.06 24.00 19.60 0.39 0.38 0.40 0.50 0.13 0.0647
+Monir 10H-10R 0.60 0.09 23.34 15.03 0.58 0.61 0.60 0.55 0.12 0.0495
+XuDudek 2H-2R 0.19 0.04 2.71 3.92 0.21 0.13 0.18 0.39 0.15 0.0822
+XuDudek 5H-5R 0.41 0.07 4.92 9.91 0.43 0.36 0.42 0.48 0.15 0.0695
+XuDudek 10H-10R 0.64 0.07 5.41 8.48 0.63 0.62 0.64 0.52 0.13 0.0535
+GuoYang 2H-2R 0.20 0.04 15.02 20.94 0.20 0.20 0.25 0.47 0.15 0.0870
+GuoYang 5H-5R 0.40 0.06 13.57 14.31 0.43 0.39 0.38 0.49 0.14 0.0677
+GuoYang 10H-10R 0.62 0.07 18.66 13.33 0.69 0.61 0.59 0.58 0.12 0.0511
+AliEtAl 2H-2R 0.19 0.04 7.57 13.29 0.18 0.20 0.24 0.38 0.15 0.0838
+AliEtAl 5H-5R 0.40 0.06 11.83 12.19 0.39 0.38 0.41 0.50 0.15 0.0660
+AliEtAl 10H-10R 0.62 0.08 11.54 9.71 0.57 0.65 0.62 0.55 0.14 0.0503
+GuoEtAl 2H-2R 0.19 0.05 11.97 17.74 0.21 0.14 0.18 0.41 0.15 0.0857
+GuoEtAl 5H-5R 0.41 0.08 14.69 13.22 0.41 0.41 0.41 0.49 0.14 0.0696
+GuoEtAl 10H-10R 0.63 0.09 17.54 13.55 0.58 0.68 0.62 0.55 0.12 0.0530
+ECT (Ours) 2H-2R 1.00 0.00 0.45 0.32 1.00 1.00 1.00 0.25 1.00 0.0321
+ECT (Ours) 5H-5R 1.00 0.00 0.40 0.20 1.00 1.00 1.00 0.32 1.00 0.0191
+ECT (Ours) 10H-10R 1.00 0.00 0.32 0.14 1.00 1.00 1.00 0.32 1.00 0.0125
+Note: SR = Success Rate, CT = Completion Time, TS = Task Success (Simple, Medium, Complex), AU = Agent Utilization, AE = Allocation Efficiency.
 IV Results and Analysis
-(a) Success rate
-(b) Average of completion time
-Figure 4: Results of the Success Rate and Average Completion Time of the Different Trust Models in team compositions.
-The star symbol indicates the best performance in each result, the highest success rate, and the shortest completion time.
-Zoom in for better readability.
-Results from our evaluation of the five trust models, including the No Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across three team compositions (e.g., 2H-2R, 5H-5R, and 10H-10R), revealed close results in both task success rates and average task completion times.
-In small team configurations (2H-2R), we observed that all trust models performed comparably in terms of success rates.
-ECT, NoTrust, and XuDudek all achieved the highest success rate of 0.35 ± 0.08, while Monir and GuoYang followed closely at 0.33 ± 0.08.
-Regarding completion times, XuDudek was the fastest at 49.28 ± 16.72 seconds, with ECT following closely at 49.66 ± 14.43 seconds.
-Notably, ECT demonstrated the lowest standard deviation in completion time, indicating more consistent performance even in small-scale operations.
-As we moved to medium-sized teams (5H-5R), the performance differences between trust models became more pronounced.
-Monir took the lead in success rate at 0.63 ± 0.10, with ECT following closely at 0.62 ± 0.10.
-The other models (NoTrust, XuDudek, and GuoYang) tied at 0.60, albeit with slightly higher standard deviations.
-Interestingly, the completion time results for this team size revealed a different pattern.
-Monir was the fastest at 48.42 ± 11.73 seconds, while ECT had the slowest mean completion time at 52.25 ± 11.25 seconds.
-This suggests a potential trade-off in ECT between maintaining a high success rate and optimizing completion time in moderately complex scenarios.
-The most notable results emerged in the large team configuration (10H-10R), where ECT demonstrated its strengths.
-In this complex scenario, ECT tied with GuoYang for the highest success rate at 0.82 ± 0.10, outperforming the other models.
-Moreover, ECT achieved the fastest mean completion time of 48.99 ± 9.13 seconds, significantly quicker than the next fastest model, Monir, at 49.14 ± 10.08 seconds.
-Also, the ECT model exhibited the lowest standard deviations in both success rate and completion time, indicating the most consistent and reliable performance in large-scale MH-MR contexts.
-An important observation from these results is the scalability demonstrated by the ECT model as its success rate improved from 0.35 in small teams to 0.82 in large teams, representing one of the most significant improvements among all models.
-This scalability suggests that the ECT model’s trust dynamics become effective as team complexity grows, a critical factor for real-world applications where scalability is often paramount.
-In the 10H-10R configuration, the ECT model outperformed other models in terms of consistency, demonstrating the lowest standard deviations in both success rate (±0.10) and completion time (±9.13s).
-V Discussion
-Non-Linearity in Team Size and Performance.
-Results from our experiment showed a non-linear relationship between team size and performance across the different trust models.
-This non-linearity aligns with findings from prior work [45] which showed that large autonomous systems often exhibit non-linear behaviors due to the complex interactions between individual robots.
-The superior performance of trust-based models, particularly the ECT model, in larger team configurations (10H-10R) suggests a potential ’scale threshold’ where trust becomes a critical factor in team efficiency and optimal performance.
-The findings from our research open up new avenues for research into the optimal team sizing and threshold where trust measurement becomes crucial for MH-MR collaborations.
-Trust as a Computational Resource.
-The ECT’s model’s performance, especially in larger teams introduces a novel perspective of trust as a form of distributed computational resource that leverages the perception of team members vis-a-vis robot performance.
-Unlike traditional views of trust as a purely psychological construct, our model and the results from the experiment suggest that trust functions as a form of collective intelligence that scales with team size and requires active participation from all team members for an objective outcome.
-This conceptualization of ’trust as a form of computation’ could help researchers reimagine how we design and implement trust models in robotic systems, potentially leading to an easily scalable and adaptable approach for eliciting and computing trust for large MH-MR teams.
-The design of the ECT model implicitly means that as the team size increases, the computational power of trust also increases.
-Task Allocation and Trust Heuristics in Large Teams.
-The performance discrepancies between trust-based and non-trust models were more pronounced in larger team configurations and raised important questions about the role of trust in mitigating complexity in MH-MR teaming and task allocation.
-As team size increases, the computational demands of task allocation grow as well.
-Our results suggest that trust might serve as an additional heuristic, in addition to other factors like fatigue and workload, to reduce this complexity by providing insight and allowing for more efficient task allocation that is based on empirical metrics in large-scale MH-MR collaboration settings.
-Specifically, the ECT model is a scalable trust model that is adaptive to different team configurations and sizes and performs relatively well regardless of team composition, showing its usefulness.
-Advantages of the ECT Model.
-Another key benefit of the ECT model lies in its direct approach.
-While other trust models often rely on independent techniques like facial detection, behavioral cues, and other indirect and after the fact approaches, the ECT model relies on direct elicitation of trust scores from human team members for a task and continuously throughout the lifecycle of the task.
-This approach ensures that the trust levels of human teams on robot performance are continuously updated.
-Furthermore, the approach of team trust aggregation allows for congregating the trust scores of all team members in measuring trust levels.
-Through this approach, the trust rating of an individual team does not have an outsized influence on the overall trust final score generated by the team.
-Study Limitations and Future Directions.
-While results from our study provided valuable insights on trust-based task allocation in MH-MR teams, it is important to also acknowledge the limitations of our study.
-First, we used simulated environments for our experiment.
-While this allows for controlled experiments that do not expose users to harm, it may not fully capture the complexities that are often present in real-world contexts.
-Furthermore, our task domain might not translate to other scenarios and domains.
-As per prior work, trust is not easy to model, and as a result, our direct question-based method might not capture all the relevant nuances required for measuring trust.
-Despite these limitations, our work provides a good foundation for future research in trust-based task allocation approaches for MH-MR teams.
-VI Conclusion
+Table II summarizes the performance of seven trust models (NoTrust, Monir, XuDudek, GuoYang, AliEtAl, GuoEtAl, and ECT) across three team configurations (2H-2R, 5H-5R, 10H-10R).
+In the 2H-2R configuration, ECT achieves a perfect success rate of 1.00 and completes tasks in just 0.45 seconds on average, far exceeding the roughly 0.19–0.20 success rates and 7.57–16.03 second completion times of the other models.
+However, ECT’s impressive speed comes with substantially greater total travel distance (639.1 versus under 50 for most competitors), resulting in a notably lower allocation efficiency (0.0321).
+Its collision rate also stands at 0.834, higher than the 0.34–0.54 range observed in other models.
+A similar pattern emerges with 5H-5R teams.
+ECT again posts a perfect success rate (1.00), compared to approximately 0.39–0.41 for the other models, and it maintains a remarkable average completion time of 0.40 seconds—an order of magnitude faster than the 11.83–27.42 seconds seen elsewhere.
+Yet ECT’s total distance climbs to 1079.6, far beyond the 100–125 range of the other approaches, while its allocation efficiency (0.0191) remains significantly lower, and its collision rate jumps to 9.363.
+In the 10H-10R configuration, ECT continues to achieve a 1.00 success rate and completes tasks in only 0.32 seconds, whereas competing models display success rates in the 0.60–0.64 range and mean completion times ranging from 5.41 to 23.59 seconds.
+This level of performance again involves a total distance of 1664.2, substantially higher than the 242–251 range for the other models, and yields an allocation efficiency of 0.0125.
+ECT’s collision rate also rises to 42.76, far surpassing the next-highest rate of around 13.64, reflecting how its continuous high-level activity drives both perfect completion and intense agent interaction.
+Overall, ECT’s trust-driven approach offers perfect success rates and extremely fast completion times for teams of all sizes.
+In scenarios where ensuring mission success and speed is paramount, such as time-critical search and rescue, these metrics may make ECT the optimal choice.
+However, the increases in total distance, collision rate, and reduced allocation efficiency highlight an inherent trade-off, as agents under ECT are more active but also more prone to collisions and resource inefficiencies.
+These results suggest that while ECT maximizes success and speed, future refinements could reduce collisions and excessive movement without compromising effectiveness.
+V Conclusion
 In this paper, we investigated the impact of trust on task allocation in multi-human multi-robot teams.
-We employed five trust models, including N-Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across different types of teams that included 2H-2R, 5H-5R, and 10H-10R.
-Our simulation task involved an SAR activity.
-Findings from our research revealed that trust-based models outperformed the no-trust model most of the time in the 10H10R team configuration and that there was no noticeable difference in the lower configurations.
-These findings have implications for trust in the context of MH-MR task allocation.
-Future work could explore trust repair mechanisms for MH-MR teams where trust has been broken due to poor task performance.
+We evaluated seven trust models, including N-Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, the Ali et al.
+Trust Model, the Guo et al.
+Trust Model, and our proposed ECT Model, across different team configurations (2H-2R, 5H-5R, and 10H-10R) in a simulated search and rescue scenario.
+Our findings indicate that the ECT model outperformed both the no-trust baseline and other trust models across most measured metrics.
+However, this performance came with a trade-off, leading to lower task allocation efficiency across all teaming formations.
+These results provide valuable insights into the role of trust in MH-MR task allocation and highlight the need for balancing efficiency with trust-based decision-making.